OpenStudy (anonymous):

i have statistics homework and im dead "Let P (A) = 0.5, P (B) = 0.7, P ( A and B ) = 0.4 Find the probability: c) Either A or B will occur d) that neither A nor B will occur if anybody is able to help, i have two more i need assistance on, but this is all i will be posting for now.

OpenStudy (amistre64):

does P(A)+P(B) = 1?

OpenStudy (amistre64):



OpenStudy (amistre64):

lets say we wanted to count X+Y, to obtain all the values X = 1+2 Y = 2+3 but X+Y is all there is, so its total should be: 1+2+3 but notice we are duplicating what they have in common, so we have to adjust for that X+Y - (XandY) gives us the total

OpenStudy (anonymous):

oh god thank you so much im v sorry for the late reply i just zoned out and tried to figure out how i was going to pass this test if i couldnt even do the prep homework thank you again